tberry1:

Given: BC = 10 inches AC = inches m∠CBD = 60° m∠CAD = 90°

Falconmaster:

Hello welcome to question cove

tberry1:

hi

Falconmaster:

hello

Falconmaster:

@Ultrilliam

tberry1:

can you help me

umm:

The area of the shaded region can be solve using the formula of the area of a circular segment A = 0.5r^2( a – sin(a)) Where r is the radius of the circle and a is the angle Solving first the area of the circular segment of the bigger circle: A = 0.5(10^2)[(60)(pi/180) – sin (60)] A = 9.0586 sq in Solving first the area of the circular segment of the smaller circle: A = 0.5(6^2)[(60)(pi/180) – sin (60)] A = 3.2611 sq in Total area = 9.0586 + 3.2611 = 12.3197 sq i

tberry1:

thanks

Ultrilliam:

Welcome to QC!

umm:

Welcome & Helo, @Ultrilliam xD

umm:

*Hello

tberry1:

A segment of a circle has a 120 arc and a chord of 8in. Find the area of the segment.

umm:

Segment 120º arc and radius of 8 inches (not chord ?) A=(120/360)∏r2 A=(1/3)(3.14159)(82) A=(1/3)(3.14159)(64) A=(1/3)(201.06176) A=67.020587 square inches